OpenStudy (anonymous):

(4/7)+(3/x-3)=53/56 x=? I need someone to show me how to do the steps plz

OpenStudy (anonymous):

I got x=11 but don't remember my exact steps

OpenStudy (rajee_sam):

yes you are right

OpenStudy (anonymous):

ok but i need the steps, cuz I forgot them

OpenStudy (rajee_sam):

First find the LCD of the two fractions on the left

OpenStudy (rajee_sam):

which is?

OpenStudy (rajee_sam):

7 ( x- 3) [4 ( x- 3 ) + 3 ( 7) ] / 7 (x-3) = 53/56 Now expand the numerator on the LHS (4x - 13 + 21) / (7x - 21) = 53/56

OpenStudy (rajee_sam):

its 12 not 13* in my previous step. 4x + 9 / 7x -21 = 53/56 Now cross multiply 56 ( 4x + 9 ) = 53 ( 7x - 21)

OpenStudy (rajee_sam):

Now again expand both sides 224 x + 504 = 371x - 1113 Combining like terms on both sides 504 + 1113 = 371x - 224x 1617 = 147x x = 1617 / 147 = 11

OpenStudy (anonymous):

ok I will try one on my own now, thanks